A stone thrown vertically up with a velocity v reaches three points A,B andC with velocities v/2,v/4,v/8 .then ab/AC=

Grade:11

1 Answers

Vikas TU
6 years ago
Dear Student,
Initial velocity = u = v
For motion upto A
Final velocity at A = v/2
Height upto A=h1
Height from A to B = h2
Height from B to C = h3
Equation of motion,
V2=U2+2As
From ground to A
=>v2/4= v2-2gh1
=>h1=3v2/8g

From A to B
=>v2/16=v2/4 – 2gh2
=>h2=3v2/32g

From B to C
=>v2/64=v2/16 – 2gh3
=>h3=3v2/128g

AB/AC = h1/h1+h2+h3
=16/21.
Cheers!!
Regards,
Vikas (B. Tech. 4th year
Thapar University)